OpenStudy (anonymous):

Use the elimination method to solve the system. x-3y=9 -x+3y=-9

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

In the first equation, you have 1x, and in the second equation, you have -1x. Since the coefficients are opposite (negatives of each other) we can eliminate it by adding the two equations together like this: x -3y = 9 -x +3y = -9 ------------ 0 + 0 = 0 0=0 Since both variables were eliminated, and you end up with a true statement, the two equations must describe the same line. So they have an infinite amount of points in common

OpenStudy (anonymous):

(5, –6) no solutions (–5, 6) infinite solutions

OpenStudy (anonymous):

choice infinite

OpenStudy (anonymous):

Thx guys

OpenStudy (anonymous):

No it will be no solution

OpenStudy (anonymous):

There will be an infinite amount of solutions, since you have two lines that lie on top of each other. They represent the same line, so every point on the line is a solution, and the line has infinitely many solutions.